OpenStudy (faman39):

Shelly delivers the weekly local paper to neighborhoods in her town. House numbers are even on one side of the street and odd on the other. Shelly delivers an equal number of papers to both sides of the street. Although she always aims for the front doorstep, Shelly typically misses on three of the tosses on her route each week. Design and conduct a simulation to estimate the probability that next week, Shelly's three misses will all be at odd-numbered houses. You can set up the experiment using 3 coins to collect the data. Allow one side of the coin to represent Heads (evens) and one

mathslover (mathslover):

what we have to do?

mathslover (mathslover):

http://answers.yahoo.com/question/index?qid=20100727190536AAGsXYG is this for which u r in the need of?

OpenStudy (anonymous):

i guess you are supposed to toss the coin three times, to represent the 3 misses then repeat the experiment by tossing another three times 3 tails would represent getting all the misses on the odd number side

OpenStudy (anonymous):

so the question comes down to, if you toss a coin three times, what is the probability that you get three tails that is not hard to answer the probability you get tails on any one toss is \(\frac{1}{2}\) and so the probability that you get tails three times in a row is \[\frac{1}{2}\times \frac{1}{2}\times \frac{1}{2}=(\frac{1}{2})^3=\frac{1}{8}\]

OpenStudy (anonymous):

of course you are supposed to set up the experiment get three coins, throw them, and see how often all land tails should be approximately one out of eight tosses you are going to have to do that yourself if you toss 3 coins many times you should see it

OpenStudy (faman39):

Thank you everyone to come here to help me! i really appreciate all of u :)

OpenStudy (faman39):

And thanks alot satellite!!!